Mid-Town Copy Service processes 1,800,000 photocopies per month at its mid-town service center. Approximately 60 percent of the photocopies require collating. Collating is currently performed by high school and college students who are paid $9 per hour. Each student collates an average of 5,000 copies per hour. Management is contemplating the lease of an automatic collating machine that has a monthly capacity of 5,000,000 photocopies, with lease and operating costs totaling $1,056, plus $0.04 per 1,000 units collated.

(a) Determine the total costs of collating 500,000 and 1,300,000 per month:

  1. With student help:

500,000

Answer

1,300,000

Answer

  1. With the collating machine:

500,000

Answer

1,300,000

Answer

(b) Determine the monthly volume at which the automatic process becomes preferable to the manual process.
Answer copies.

(c) Should Mid-Town Copy lease the automatic collating machine at this time?

Given the current cost structure Midtown should not lease the collating machine regardless of the monthly volume.

Midtown should lease the collating machine at the current monthly volume.

Midtown should lease the collating machine only if the monthly volume is less than 600,000 copies.

Midtown should only lease the collating machine if the monthly volume is more than 500,000 copies.
Answer #1

a.

With student help

Collating 500,000 = $ 9 * 500,000 / 5,000 = $ 900

Collating 1,300,000 = $ 9 * 1,300,000 / 5,000 = $ 2,340

With the collating machine

Collating 500,000 = $ 1,056 + 0.04 * 500,000/1000 = 1056 +20 = $ 1076

Collating 1,300,000 = $ 1,056 + 0.04 * 1300,000/1000 = 1056 + 52 = $ 1108

b.

$1056+$0.04*X/1,000=$9*X/5,000

Or, 1056 + 0.00004x = 0.0018x

Or, x = 1056/0.00176 = 600,000

When the monthly volume is over 600,000, the automatic process becomes preferable to the manual process.

c. Midtown should lease the collating machine at the current monthly volume.

Current volume = 1800,000 * 60 % = 1,080,000

If collating machine is used, the cost will be less than that if student help is required.

Since, the volume is preferable over 600,000, the automatic process is recommended and midtown should lease collating machine at current monthly volume

kindly upvote